A 10-year $1,000 bond sells for $980, and the flotation costs are 2% of the par value. The coupon rate is 8%. Put in the values of the variables and calculate the before-tax cost of capital from this bond?

Coupon dollars (I) =   

Flotation costs in dollars (FC) =   

term (n) =   

Price (P) =   

Net proceeds (N_p) =   

rd =   %
Answer #1
As the question is silent, on whether the bond pays annual or semi-annual coupons,
we shall calculate for both cases .
Before-tax cost of the bond=
Net proceeds=(( $ Coupon)*(1-(1+r)^-n)/r)+(FV/(1+r)^n)
Where,
Net proceeds= Selling price-Flotation cost,ie. 980-(1000*2%)= 960
Coupon amt.= (1000*8%)= $ 80 , if annual or   80/2= $ 40 ,if semi-annual coupon
n= no.of coupon periods, ie. 10 for annual & 10*2= 20 for semi-annual coupon bonds
So, substituting the values,
For annual coupon bonds
960=(80*(1-(1+r)^-10)/r)+(1000/(1+r)^10
Solving for r, we get the annual before-tax cost, r as:
8.61%
For semi-annual coupon bonds
960=(40*(1-(1+r)^-20)/r)+(1000/(1+r)^20
Solving for r, we get the annual before-tax semi-annual cost, r as:
4.3023%
Converting,
The annual before tax cost =
(1+4.3023%)^2-1=
8.79%
Before-tax cost of capital of this bond is its yield that equals the present net proceeds , to the present value of its future--- coupon +principal-at-maturity -- cash flows
Coupons
Annual Semi-annual
Coupon dollars (I) =1000*8%= 80 40
Flotation costs in dollars (FC) =(1000*2%) = 20 20
term (n) =    10 10 10
Price (P) = 980   980 980
Net proceeds (N_p) =   980-20= 960 960
rd =   %   semi-annual                                    4.30%
                  Annual 8.61% 8.79%